OpenStudy (anonymous):

integrate 1/(xln(x^3))

OpenStudy (anonymous):

substitute: u=lnx, so du=1/x so, your int becomes: int(1/u^3)du=-1/2 * (1/u^2)=-1/2 * 1/(ln x) +const